OpenStudy (sissyedgar):

ASAP HELP PLEASE!!! MEDALS! Which equation demonstrates the opposite of a sum property? A. (22 + 10) + 3 = (10 + 22) + 3 B. –5 + 13 = 13 + (–5) C. (12 + (–4)) + 14 = 12 + ((–4) + 14) D. –(–6 + 17) = 6 + (–17)

OpenStudy (sissyedgar):

@cheerleader

OpenStudy (cheerleader):

a

OpenStudy (sissyedgar):

thx bb

OpenStudy (cheerleader):

yeah no prob bb

OpenStudy (sissyedgar):

i didnt like how the other 1 looked... –(z + (–17)) = 12 + 17 A. –46 B. –12 C. 12 D. 46

OpenStudy (sissyedgar):

darn ???

OpenStudy (cheerleader):

c

OpenStudy (sissyedgar):

thx

OpenStudy (sissyedgar):

duh me! lol

OpenStudy (sissyedgar):

What is the value of the expression? –(5 + (–11.3)) + (–4.3) A. –20.6 B. –10.6 C. –2 D. 2

OpenStudy (cheerleader):

a

OpenStudy (sissyedgar):

muah

OpenStudy (cheerleader):

muah

OpenStudy (sissyedgar):

i MIGHT have 1 or 2 more

OpenStudy (sissyedgar):

What value of c makes the statement true? c – (–17.6) = –4 A. –21.6 B. –12.4 C. 12.4 D. 21.6

OpenStudy (cheerleader):

is this adding or multiplying?

OpenStudy (sissyedgar):

idk, this is al i got from the question

OpenStudy (sissyedgar):

nvm i got it, addition, A)

OpenStudy (sissyedgar):

last 1... The temperature at noon was –6.5⁰C. The temperature rose 3.5⁰C from noon to 1 p.m. The temperature increased t degrees centigrade from 1 p.m. to 2 p.m. The temperature at 2 p.m. was –12⁰C. Which equation can be used to solve for t? A. –6.5 – 3.5 – t = –12 B. –6.5 + 3.5 + t = –12 C. –6.5 + 3.5 – t = –12 D. –6.5 – 3.5 + t = –12

OpenStudy (sissyedgar):

?????, uuuuuhhhhhgggggg

OpenStudy (cheerleader):

i think it c

OpenStudy (sissyedgar):

ok, thx soooo much hun, i cant put thw other 1 up cause ist not words its pics so ill have 2 guess

OpenStudy (cheerleader):

ok bb

OpenStudy (cheerleader):

anything else babe